# Quasi-perfect blowups detect regularity

### v1

@@ 2.3. Morphisms
  • a modification if it is proper and birational (i.e. generically an isomorphism), see e.g. [Sta25, Tag 0ACU] for some characterizations and
  • an alteration if it is proper and generically finite, see e.g. [Sta25, Tag 0AD7] for some characterizations.
Usually, in this work the target (and hence, source) of the morphism will be Noetherian. Note that, for either type, the morphism above is, in fact, surjective. This follows as they are by definition dominant (this is implicit in generically finite/an isomorphism) and a dominant proper morphism is surjective. The modifications we will be interested in below arise as blowups. See [Sta25, Tag 085P] some background on blowups of algebraic spaces.
@@
3. Results
To start, we note an ‘obvious’ lemma which is recorded for convenience.
@@ 3. Results
Let $f\colon Y\to X$ be a morphisms of quasi-compact quasi-separated algebraic spaces with $Y$ nonempty. Then $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is nonzero; in particular, $\mathbf{R}f_{\ast}G$ is nonzero for any compact generator $G$ of $D_{\operatorname{qc}}(Y)$.
@@
The second claim follows from the first as $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is finitely built from $\mathbf{R}f_{\ast}G$ (i.e. $\mathcal{O}_{Y}\in\langle G\rangle$ and so also after applying $\mathbf{R}f_{\ast}$). To see the first claim, observe that it follows from pullback/pushforward adjunction that
$0$ $\not=H^{0}(Y,\mathcal{O}_{Y})$
$\cong\operatorname{Ext}^{0}(\mathcal{O}_{Y},\mathcal{O}_{Y})$
$\cong\operatorname{Ext}^{0}(\mathbf{L}f^{\ast}\mathcal{O}_{X},\mathcal{O}_{Y})$
$\cong\operatorname{Ext}^{0}(\mathcal{O}_{X},\mathbf{R}f_{\ast}\mathcal{O}_{Y})$
and so $\mathbf{R}f_{\ast}\mathcal{O}_{Y}$ is nonzero. ∎
The next should be compared to [Sta25, Tag 00OC] and gives a ‘spacy’ version of this.
@@
Let $X$ be a decent (e.g. a Noetherian) algebraic space. Suppose $x\in|X|$ is a closed point such that $i_{x,*}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}$ is perfect, where $i_{x}\colon{\operatorname{Spec}}(\kappa(x))\to X$ denotes the residue field of $X$ at $x$. Then $x$ is a regular point of $X$.
@@
Let $(U,u)\to(X,x)$ be the ‘elementary étale neighborhood’ of [Sta25, Tag 0BBP], i.e. $U$ is an affine scheme, $u$ is the only point of $U$ lying over $x$, and the induced homomorphism of residue fields $\kappa(x)\to\kappa(u)$ is an isomorphism. We claim that the pullback of $x$ under $f\colon U\to X$ can be identified with $u$, i.e. we have a cartesian square
@@
Indeed, let $U_{x}$ denote the pullback. It follows from categorical nonsense that $U_{x}\to U$ is a monomorphism (as $\operatorname{Spec}(\kappa(x))\to X$ is). Moreover, as $U_{x}\to\operatorname{Spec}\kappa(x)$ is étale, it is a disjoint union of fields. However, as each field corresponds to a distinct point lying over $x$ (they all go to different points as monomorphisms of schemes are injective), $U_{x}$ is the spectrum of a field. Consequently, we must have $U_{x}=\operatorname{Spec}(\kappa(u))$ as $U_{x}\to U$ is a monomorphism. In particular, it also follows that $u$ is necessarily a closed point as $i_{x}$ is a closed immersion by e.g. [Sta25, Tag 0H1R & Tag 0H1U].
Next, using e.g. flat base change, one sees
$f^{*}i_{x,*}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}=i_{u,*}\mathcal{O}_{\operatorname{Spec(\kappa(u))}}$
and so also perfect. Taking the stalk at $u$, one has the residue field $\kappa(u)$ is a perfect $\mathcal{O}_{U,u}$-module. Thus, $u$ is a regular point by [Sta25, Tag 00OC] and so $x$ is a regular point of $X$ by definition. ∎
Theorem 3.3.
A Noetherian algebraic space $X$ is regular if, and only if, the blowup of $X$ along any closed point is a quasi-perfect morphism.
Proof.
If $X$ is regular, then for any proper $f$ we have
Hence, any blowup is quasi-perfect as these are proper.
For the converse, it suffices to show that any closed point is regular (see e.g. [DLMRP25, Lemma 3.2]). So, let $p\in|X|$ be a closed point and denote by $i\colon\operatorname{Spec}(\kappa(p))\to X$ its associated residue field. As the point is closed, $i$ is a closed immersion; see e.g. [Sta25, Tag 0H1R & Tag 0H1U]. Consider the fibered square
[…]
where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor. The hypothesis tells us that $f$ is a quasi-perfect morphism.
As the exceptional divisor $E$ is an effective Cartier divisor on $X^{\prime}$, we have a short exact sequence
$0\to\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to 0$
from which we obtain a distinguished triangle in $D^{b}_{\operatorname{coh}}(X^{\prime})$,
(3.2) $\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to\mathcal{O}_{X^{\prime}}(-E)[1].$
Furthermore, as $\mathcal{O}_{X^{\prime}}(-E)$ is an invertible sheaf, and so perfect, we have $j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X^{\prime})$.
Hence, as $f$ is quasi-perfect by assumption, we know that $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$. Moreover, from the commutativity of Equation 3.1 and Lemma 3.1, we see that
@@
It follows that $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect. Hence $p$ is regular by Lemma 3.2 and we are done. ∎
Remark 3.4.
It is worthwhile to stress the following: the blowup at a closed point being quasi-perfect implies that that point is regular. Indeed, this is essentially the content of the above proof.
@@
The blowup is reduced, see e.g. [Sta25, Tag 085W], as the latter is an étale local property. Moreover, since blowups are representable, it follows that $X^{\prime}$ is decent (see e.g. [Sta25, Tag 0ABT]).
Thus, it remains to show that $|X^{\prime}|$ is irreducible, which is more subtle as this notion is not étale local. Let $Z$ be the center of the blow-up. We have that $f^{-1}(Z)$ is an effective Cartier divisor on $X^{\prime}$ and that $f$ induces an isomorphism $f^{-1}(X\setminus Z)=X\setminus Z$; see e.g. [Sta25, Tag 085T]. When $|X|$ is irreducible so is the open subset $|X\setminus Z|$ and so $|f^{-1}(X\setminus Z)|=|X^{\prime}\setminus f^{-1}(Z)|$ is also. However, $f^{-1}(Z)$ being effective Cartier implies that the open subspace $X^{\prime}\setminus f^{-1}(Z)$ is scheme-theoretically dense in $X^{\prime}$, see [Sta25, Tag 083S]. In particular, the closure of $|X^{\prime}\setminus f^{-1}(Z)|$ (in $|X^{\prime}|$) is $|X^{\prime}|$ and so the latter is also irreducible.
To see that $f$ is a modification when the ideal is of finite type, note that in this case $f$ is proper by [Sta25, Tag 085Z]. ∎
@@
An integral Noetherian algebraic space $X$ is regular if, and only if, every alteration (resp. modification) of $X$ is quasi-perfect.
@@
This follows immediately from Theorem 3.3. Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma 3.5. Moreover, as modifications are alterations the case of alterations also follows. ∎
@@
In the special case where there exists an alteration or modification $f\colon Y\to X$ of algebraic spaces where $Y$ is regular and $f$ quasi-perfect, then one can show $X$ regular more easily using ‘proper descent’. For example, this follows from [HLLP25, Corollary 6.3], and the fact that $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$ implies regularity. So, the novelty of Corollary 3.6 lies in its independence of the existence of such a morphism, giving a different characterization of regularity.
Next come some lemmas that allow one to check quasi-perfectness using flat covers.
@@
A morphism $f\colon Y\to X$ of quasi-compact quasi-separated algebraic spaces is quasi-perfect if, and only if, $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$ for any compact generator $P$ of $Y$.
@@
    If $f$ is quasi-perfect and $s$ is affine, then $f^{\prime}$ is quasi-perfect.
@@
To show the first claim, first note that when $s$ is an affine morphism so is its base change $s^{\prime}$. Consequently, $\mathbf{R}s^{\prime}_{\ast}=s^{\prime}_{\ast}$ is conservative (this statement is local on the target and so reduces to the case of schemes) and it follows that for any compact generator $Q$ of $Y$ its pullback $(s^{\prime})^{\ast}Q$ is a compact generator of $Y\times_{X}U$ (see e.g. [Sta25, Tag 0E4R]). Thus, as $f$ is assumed quasi-perfect, $\mathbf{R}f_{\ast}Q$ is perfect and so also is its pullback $s^{\ast}\mathbf{R}f_{\ast}Q$. From flat base change, it follows that
@@
Hence, Lemma 3.8 implies $f^{\prime}$ is a quasi-perfect morphism.
For the second claim, assume now that $f$ is surjective and $f^{\prime}$ is quasi-perfect. Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation 3.4. As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact. Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma 2.1. By Lemma 3.8, we have that $f$ is quasi-perfect, which completes the proof. ∎
@@
The forward direction follows from Lemma 3.9. For the converse, denote the base change of $f$ along $s$ as in Equation 3.3. Again, choose a compact generator $Q$ of $Y$. By assumption $f^{\prime}$ is quasi-perfect, hence, using Equation 3.4, $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect. However, as $s$ corresponds to a faithfully flat ring map, it follows from [Sta25, Tag 068T] that $\mathbf{R}f_{\ast}Q$ is perfect. Thus, Lemma 3.8 shows that $f$ is quasi-perfect. ∎
Using Lemma 3.9, we obtain an alternative proof of the fact that regularity can be checked fpqc locally.
@@
Let $X$ be a separated Noetherian algebraic space. If there is an fpqc cover $\{f_{i}\colon U_{i}\to X\}_{i\in I}$ where each $U_{i}$ is regular, then $X$ is regular.
@@
First, as $X$ is quasi-compact, we can reduce to an fpqc cover with finitely many components (i.e. $|I|<\infty$). Secondly, if needed, we can refine the cover so that each $U_{i}$ is an affine scheme (e.g. take étale surjective morphisms from affine schemes). Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is affine by [Sta25, Tag 08GB] as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme. To show $X$ is regular, it suffices by Theorem 3.3 to show that blowups along nonzero ideal sheaves are quasi-perfect. However, as blowups commute with flat base change, see e.g. [Sta25, Tag 0805], we have that the base change to $U$ is again such a blow-up. Consequently, as $U$ is regular, it is quasi-perfect. Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma 3.9, showing the claim. ∎
@@
There are three important cases of faithfully flat morphisms that we will apply Lemma 3.10 to.
@@
There are local ring morphisms $R\to\widehat{R}$ and $R\to R^{h}\to R^{sh}$ which are flat and consequently faithfully flat; see e.g. [Sta25, Tag 00MC and Tag 07QM]. (As an aside, there is also a flat local ring morphism $R^{h}\to\widehat{R}$, but we do not need it).
@@
Let $f\colon Y\to X$ be a proper morphism of algebraic spaces with $X$ separated and Noetherian. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
@@
The proof of ($\star$3) is similar to that of Lemma 3.9; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work. Pick a point $p\in U$ and look at the fibered square
@@
Again, choose any compact generator $Q$ of $Y$. The hypothesis implies that $\mathbf{R}(f_{p})_{\ast}t^{\ast}_{p}Q$ is perfect, and so flat base change tells us $s_{p}^{\ast}\mathbf{R}f_{\ast}Q=(s^{\ast}\mathbf{R}f_{\ast}Q)_{p}$ is perfect. As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper. Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma 2.1 and $f$ is quasi-perfect by Lemma 3.8. ∎
@@
Let $f\colon Y\to X$ be a proper morphism of algebraic spaces with $X$ separated and Noetherian and let $A_{\#}$ and $s\colon U\to X$ be as in Theorem 3.14. Then the locus
(3.5) $S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
@@
By the proof of Theorem 3.14 (i.e. using Lemma 3.10) we know that
$S_{X}:=$ $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{sh}\textrm{ is quasi-perfect}\right\}$
@@
$S_{U}:=$ $\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}$
@@
To this end, observe by Equation 2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$. As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma 3.10. Therefore,
@@
As a consequence of Equation 3.6 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open. Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
@@
One can think of the subset $S$ appearing in Corollary 3.15 as the ‘quasi-perfect locus of $f$’. It follows from the first part of the proof that in general, i.e. regardless of properness of $f$, one can compute it étale locally (Equation 3.6) and that the specific choice of $A_{\#}$ does not matter. Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
@@
Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [Sta25, Tag 04NA] for the relevant definition. whose étale local ring is reduced (e.g. this is the case when $X$ is reduced). Suppose further that $f\colon Y\to X$ is a proper surjective morphism. Then the regular locus of $X$ contains a nonempty open subset (i.e. is JJ-0) whenever $Y$ has this property.
@@
  • As $X$ admits an open dense subscheme, which in particular contains $\eta$, see e.g. [Sta25, Tag 06NH], we may assume that $X$ is a scheme. Moreover, by choosing any affine open neighborhood $U$ of $\eta$ and by restricting to $f^{-1}(U)\to U$, we may assume $X$ is affine and, in particular, separated.
  • Next, let $S$ denote the open subset from Corollary 3.15; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty. It follows from Theorem 3.14 that the restriction $f^{-1}(S)\to S$ is quasi-perfect. So we may assume $f$ is quasi-perfect.
@@
In order to finish the proof, is suffices to show that $X$ is regular. This follows from ‘proper descent’, as already indicated in Remark 3.7; but we give a self-contained proof, more in line with that of Theorem 3.3, for the convenience of the reader.

### v2

@@ 2.3. Morphisms
  • a modification if it is proper and birational (i.e. generically an isomorphism), see e.g. [STA26, Tag 0ACU] for some characterizations and
  • an alteration if it is proper and generically finite, see e.g. [STA26, Tag 0AD7] for some characterizations.
Usually, in this work the target (and hence, source) of the morphism will be Noetherian. Note that, for either type, the morphism above is, in fact, surjective. This follows as they are by definition dominant (this is implicit in generically finite/an isomorphism) and a dominant proper morphism is surjective. The modifications we will be interested in below arise as blowups. See [STA26, Tag 085P] for some background on blowups of algebraic spaces.
@@
3. Quasi-perfect blowups detect regularity
This section shows a relationship between regularity and having ‘enough’ quasi-perfect ‘probing morphisms’. The following should be compared to [STA26, Tag 00OC] and gives a ‘spacy’ version of this.
@@ 3. Quasi-perfect blowups detect regularity
Let $X$ be a decent (e.g. a Noetherian) algebraic space. Suppose $x\in|X|$ is a closed point such that $i_{x,\ast}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}$ is perfect, where $i_{x}\colon{\operatorname{Spec}}(\kappa(x))\to X$ denotes the residue field of $X$ at $x$. Then $x$ is a regular point of $X$.
@@
Let $(U,u)\to(X,x)$ be the ‘elementary étale neighborhood’ of [STA26, Tag 0BBP], i.e. $U$ is an affine scheme, $u$ is the only point of $U$ lying over $x$, and the induced homomorphism of residue fields $\kappa(x)\to\kappa(u)$ is an isomorphism. We claim that the pullback of $x$ under $f\colon U\to X$ can be identified with $u$, i.e. we have a cartesian square
[…]
Indeed, let $U_{x}$ denote the pullback. It follows from categorical nonsense that $U_{x}\to U$ is a monomorphism (as $\operatorname{Spec}(\kappa(x))\to X$ is). Moreover, as $U_{x}\to\operatorname{Spec}(\kappa(x))$ is étale, it is a disjoint union of fields. However, as each field corresponds to a distinct point lying over $x$ (they all go to different points as monomorphisms of schemes are injective), $U_{x}$ is the spectrum of a field. Consequently, we must have $U_{x}=\operatorname{Spec}(\kappa(u))$ as $U_{x}\to U$ is a monomorphism. In particular, it also follows that $u$ is necessarily a closed point as $i_{x}$ is a closed immersion by e.g. [STA26, Tag 0H1R & Tag 0H1U].
Next, using flat base change, one sees
$f^{\ast}i_{x,\ast}\mathcal{O}_{\operatorname{Spec(\kappa(x))}}=i_{u,\ast}\mathcal{O}_{\operatorname{Spec(\kappa(u))}}$
and so also perfect. Taking the stalk at $u$, one has the residue field $\kappa(u)$ is a perfect $\mathcal{O}_{U,u}$-module. Thus, $u$ is a regular point by [STA26, Tag 00OC] and so $x$ is a regular point of $X$ by definition. ∎
@@
Let $X$ be Noetherian algebraic space $X$. If the blowup of $X$ along a closed point $p\in|X|$ induces a quasi-perfect morphisms, then $p$ is regular.
@@
Denote by $i\colon\operatorname{Spec}(\kappa(p))\to X$ the residue field associated to our closed point $p$; this is a closed immersion, see e.g. [STA26, Tag 0H1R & Tag 0H1U]. Consider the fibered square
@@
where $f$ is the blowup of $X$ at $p$ and $E$ is the exceptional divisor. By definition of blowing up the exceptional divisor $E$ is an effective Cartier divisor, and so it follows from the distinguished triangle
$\mathcal{O}_{X^{\prime}}(-E)\to\mathcal{O}_{X^{\prime}}\to j_{\ast}\mathcal{O}_{E}\to\mathcal{O}_{X^{\prime}}(-E)[1].$
that $j_{\ast}\mathcal{O}_{E}$ is perfect. Hence, as $f$ is quasi-perfect by assumption, $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}$ is perfect too. To conclude, observe from the commutativity of Diagram˜3.1 that
@@
Hence $i_{\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p))}$ is a summand of $\mathbf{R}f_{\ast}j_{\ast}\mathcal{O}_{E}\in\operatorname{Perf}(X)$ and so is also perfect. Thus, $p$ is regular by Lemma˜3.1, and we are done. ∎
Proposition 3.3.
A Noetherian algebraic space $X$ is regular if and only if the blowup of $X$ along any closed point is a quasi-perfect morphism.
Proof.
If $X$ is regular, then for any proper $f$ we have
Hence, any blowup is quasi-perfect as these are proper. The converse follows from Lemma˜3.2 as it suffices to check regularity at closed points (see e.g. [DLR+25, Lemma 3.2]). ∎
@@
The blowup is reduced, see e.g. [STA26, Tag 085W], as the latter is an étale local property. Moreover, since blowups are representable, it follows that $X^{\prime}$ is decent (see e.g. [STA26, Tag 0ABT]).
Thus, it remains to show that $|X^{\prime}|$ is irreducible, which is more subtle as this notion is not étale local. Let $Z$ be the center of the blow-up. We have that $f^{-1}(Z)$ is an effective Cartier divisor on $X^{\prime}$ and that $f$ induces an isomorphism $f^{-1}(X\setminus Z)=X\setminus Z$; see e.g. [STA26, Tag 085T]. When $|X|$ is irreducible so is the open subset $|X\setminus Z|$ and so $|f^{-1}(X\setminus Z)|=|X^{\prime}\setminus f^{-1}(Z)|$ is also. However, $f^{-1}(Z)$ being effective Cartier implies that the open subspace $X^{\prime}\setminus f^{-1}(Z)$ is scheme-theoretically dense in $X^{\prime}$, see [STA26, Tag 083S]. In particular, the closure of $|X^{\prime}\setminus f^{-1}(Z)|$ (in $|X^{\prime}|$) is $|X^{\prime}|$ and so the latter is also irreducible.
To see that $f$ is a modification when the ideal is of finite type, note that in this case $f$ is proper by [STA26, Tag 085Z]. ∎
@@
An integral Noetherian algebraic space $X$ is regular if and only if every alteration (resp. modification) of $X$ is quasi-perfect.
@@
This follows immediately from Lemma˜3.2. Indeed, the blowups appearing in the proof have integral target and so are modifications by Lemma˜3.4. Moreover, as modifications are alterations the case of alterations also follows. ∎
@@
In the special case where there exists an alteration or modification $f\colon Y\to X$ of algebraic spaces where $Y$ is regular and $f$ quasi-perfect, then one can show $X$ regular more easily using ‘proper descent’. For example, this follows from [HLL+25, Corollary 6.3], and the fact that $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$ implies regularity. So, the novelty of Corollary˜3.5 lies in its independence of the existence of such a morphism, giving a different characterization of regularity.
4. On proper quasi-perfect morphisms
Here we study quasi-perfect morphisms which are proper. We start with a few lemmas that allow one to check quasi-perfectness using flat covers.
@@ 4. On proper quasi-perfect morphisms
A morphism $f\colon Y\to X$ of quasi-compact quasi-separated algebraic spaces is quasi-perfect if and only if $\mathbf{R}f_{\ast}P\in\operatorname{Perf}(X)$ for any compact generator $P$ of $Y$.
@@
    If $f$ is quasi-perfect and $s$ is quasi-affine, then $f^{\prime}$ is quasi-perfect.
@@
To show the first claim, first note that when $s$ is an quasi-affine morphism so is its base change $s^{\prime}$. Consequently, $\mathbf{R}s^{\prime}_{\ast}$ is conservative (see e.g. [HR17, Corollary 2.8]) and it follows that for any compact generator $Q$ of $Y$ its pullback $(s^{\prime})^{\ast}Q$ is a compact generator of $Y\times_{X}U$ (see e.g. [STA26, Tag 0E4R]). Thus, as $f$ is assumed quasi-perfect, $\mathbf{R}f_{\ast}Q$ is perfect and so also is its pullback $s^{\ast}\mathbf{R}f_{\ast}Q$. From flat base change, it follows that
@@
Hence, Lemma˜4.1 implies $f^{\prime}$ is a quasi-perfect morphism.
For the second claim, assume now that $f$ is surjective and $f^{\prime}$ is quasi-perfect. Then $\mathbf{R}f^{\prime}_{\ast}(s^{\prime})^{\ast}Q$ is perfect and so is $s^{\ast}\mathbf{R}f_{\ast}Q$ by Equation˜4.2. As $s$ is surjective, so is $s^{\prime}$, and so it is an fpqc cover as it is quasi-compact. Therefore, $\mathbf{R}f_{\ast}Q$ is perfect as this can be checked flat locally by Lemma˜2.1. By Lemma˜4.1, we have that $f$ is quasi-perfect, which completes the proof. ∎
@@
The forward direction follows from Lemma˜4.2. For the converse, denote the base change of $f$ along $s$ as in Diagram˜4.1. Again, choose a compact generator $Q$ of $Y$. By assumption $f^{\prime}$ is quasi-perfect, hence, using Equation˜4.2, $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect. However, as $s$ corresponds to a faithfully flat ring map, it follows from [STA26, Tag 068T] that $\mathbf{R}f_{\ast}Q$ is perfect. Thus, Lemma˜4.1 shows that $f$ is quasi-perfect. ∎
Using Lemma˜4.2, we obtain an alternative proof of the fact that regularity can be checked fpqc locally. Before doing so, we recall a fact whose proof we add for lack of finding a reference.
Lemma 4.4.
Let $X$ be a quasi-separated algebraic space. Then any morphism from an affine scheme to $X$ is quasi-affine.
Proof.
Suppose $U\to X$ is a map from an affine scheme. So show it is affine, by definition, we need to show, for any affine scheme $V$ and morphisms $V\to X$, that $U\times_{X}V$ is quasi-affine. As the latter is just the pullback of the morphisms $U\times V\to X\times X$ along the diagonal $X\to X\times X$, this follows from the fact the diagonal of $X$ is quasi-affine, see e.g. [STA26, Tag 03HK & Tag 02LR]. ∎
@@
Let $X$ be a Noetherian algebraic space. If there is an fpqc cover $\{f_{i}\colon U_{i}\to X\}_{i\in I}$ where each $U_{i}$ is regular, then $X$ is regular.
@@
First, as $X$ is quasi-compact, we can reduce to an fpqc cover with finitely many components (i.e. $|I|<\infty$). Secondly, if needed, we can refine the cover so that each $U_{i}$ is an affine scheme (e.g. take étale surjective morphisms from affine schemes). Next, consider the induced morphism $f\colon\bigsqcup_{i\in I}U_{i}\to X$ which is faithfully flat by construction and moreover is quasi-affine by Lemma˜4.4 as $U:=\bigsqcup_{i\in I}U_{i}$ is (isomorphic to) an affine scheme. To show $X$ is regular, it suffices by Lemma˜3.2 to show that blowups along nonzero ideal sheaves are quasi-perfect. However, as blowups commute with flat base change, see e.g. [STA26, Tag 0805], we have that the base change to $U$ is again such a blow-up. Consequently, as $U$ is regular, it is quasi-perfect. Thus, the original blow-up of $X$ is also quasi-perfect, by Lemma˜4.2, showing the claim. ∎
@@
There are three important cases of faithfully flat morphisms that we will apply Lemma˜4.3.
@@
There are local ring morphisms $R\to\widehat{R}$ and $R\to R^{h}\to R^{sh}$ which are flat and consequently faithfully flat; see e.g. [STA26, Tag 00MC and Tag 07QM]. (As an aside, there is also a flat local ring morphism $R^{h}\to\widehat{R}$, but we do not need it).
@@
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose $A_{\#}\in\{\mathcal{O}_{X,\#}^{h},\mathcal{O}_{X,\#}^{sh}\}$. Then the following are equivalent:
@@
The proof of ($\star$4) is similar to that of Lemma˜4.2; however, note that $\operatorname{Spec}(\mathcal{O}_{U,p})\to X$ is generally not surjective so we need to do some extra work. Pick a point $p\in U$ and look at the fibered square
@@
Again, choose any compact generator $Q$ of $Y$. The hypothesis implies that $\mathbf{R}(f_{p})_{\ast}t^{\ast}_{p}Q$ is perfect, and so flat base change tells us $s_{p}^{\ast}\mathbf{R}f_{\ast}Q=(s^{\ast}\mathbf{R}f_{\ast}Q)_{p}$ is perfect. As $U$ is Noetherian, and as $p\in U$ was arbitrary, we have that $s^{\ast}\mathbf{R}f_{\ast}Q$ is perfect (this can be checked affine locally via [AIL10, Theorem 4.1])–this uses the fact that $s^{\ast}\mathbf{R}f_{\ast}Q$ is bounded and coherent by virtue of $f$ being proper. Consequently, so is $\mathbf{R}f_{\ast}Q$ by Lemma˜2.1 and $f$ is quasi-perfect by Lemma˜4.1. ∎
@@
Let $f\colon Y\to X$ be a proper morphism of Noetherian algebraic spaces. Suppose that $A_{\#}$ and $s\colon U\to X$ are as in Theorem˜4.8. Then the locus
$S:=\left\{p\mid\text{the projection }Y\times_{X}\operatorname{Spec}(A_{p})\to\operatorname{Spec}(A_{p})\textrm{ is quasi-perfect}\right\}$
@@
By the proof of Theorem˜4.8 (i.e. using Lemma˜4.3) we know that
$S_{X}$= $\left\{x\in|X|\mid\text{base change to }\mathcal{O}_{X,x}^{sh}\textrm{ is quasi-perfect}\right\}$
@@
$S_{U}$= $\left\{u\in U\mid\text{base change to }\mathcal{O}_{U,u}\textrm{ is quasi-perfect}\right\}$
@@
To this end, observe by Equation˜2.1 that $\mathcal{O}_{X,s(u)}^{sh}=\mathcal{O}^{sh}_{U,u}$ for any $u\in U$. As $\mathcal{O}_{U,u}\to\mathcal{O}^{sh}_{U,u}$ is faithfully flat, quasi-perfectness of base change to either of them is equivalent by Lemma˜4.3. Therefore,
@@
As a consequence of Equation˜4.3 and that $s\colon U\to|X|$ is an open quotient map, it follows that $S_{X}$ is open if and only if $S_{U}$ is open. Thus, picking another cover if necessary we may assume $U$ is affine and it suffices to show $S_{U}$ is open to finish the proof.
@@
One can think of the subset $S$ appearing in Corollary˜4.9 as the ‘quasi-perfect locus of $f$’. It follows from the first part of the proof that in general, i.e. regardless of properness of $f$, one can compute it étale locally (Equation˜4.3) and that the specific choice of $A_{\#}$ does not matter. Moreover, this subset is open when $f$ is proper; it would be interesting to know whether or not this holds for more general morphisms.
@@
Let $X$ be a Noetherian algebraic space containing at least one codimension 0 point [^1] See [STA26, Tag 04NA] for the relevant definition. whose étale local ring is reduced (e.g. this is the case when $X$ is reduced). Suppose further that $f\colon Y\to X$ is a proper surjective morphism. Then the regular locus of $X$ contains a nonempty open subset (i.e. is JJ-0) whenever $Y$ has this property.
@@
  • As $X$ admits an open dense subscheme, which in particular contains $\eta$, see e.g. [STA26, Tag 06NH], we may assume that $X$ is a scheme. Moreover, by choosing any affine open neighborhood $U$ of $\eta$ and by restricting to $f^{-1}(U)\to U$, we may assume $X$ is affine and, in particular, separated.
  • Next, let $S$ denote the open subset from Corollary˜4.9; as $\eta\in S$, since $\mathcal{O}_{X,\eta}$ is a field by assumption, $S$ is non-empty. It follows from Theorem˜4.8 that the restriction $f^{-1}(S)\to S$ is quasi-perfect. So we may assume $f$ is quasi-perfect.
@@
In order to finish the proof, is suffices to show that $X$ is regular. This follows from ‘proper descent’, as already indicated in Remark˜3.6; but we give a self-contained proof, more in line with that of Lemma˜3.2, for the convenience of the reader.